30% of my product is priced at $4.21
40% of my product is priced at $19.15
15% of my product is priced at $38.30
15% of my product is priced at $0.00
What is my average price?

Answer by checkley71(8403)

.3*4.21+.4*19.15+.15*38.30+.15*0=1.263+7.66+5.745+0=14.668 is the average price of all products.